Grade 12Differential Calculus

Differential of log^n x

milind
12 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
f(x) = log^{n}(x)
f'(x) = n.log^{n-x}(x).\frac{1}{x}
f'(x) = \lim_{h\rightarrow 0}\frac{f(x+h)-f(x)}{h}
f'(x) = \lim_{h\rightarrow 0}\frac{(log(x+h))^{n}-(log(x))^{n}}{h}
It is a zero by zero form, so apply L’Hospital rule
f'(x) = \lim_{h\rightarrow 0}\frac{n(log(x+h))^{n-1}}{1}.\frac{1}{x+h}
f'(x) = n(log(x))^{n-1}.\frac{1}{x}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty